OpenStudy (anonymous):

Which expression represents in radical form 3x^5/6?

OpenStudy (anonymous):

seems like question is incomplete

OpenStudy (anonymous):

3 ^6sqrtx5 x ^6sqrtx5 32x^5 16x^6

OpenStudy (anonymous):

those are the expressions woops

OpenStudy (dumbcow):

\[3x^{5/6} = 3\sqrt[6]{x^{5}}\]

jimthompson5910 (jim_thompson5910):

\[\large 3x^{\frac{5}{6}}=3\sqrt[6]{x^5}\]

OpenStudy (dumbcow):

denominator of rational exponent is always the root